OpenStudy (anonymous):

Hi! I have a project on applications of rational equations. I've done 75% of the work, but need help walking through the problem I've created.

OpenStudy (phebe):

ok

OpenStudy (anonymous):

I take twice as long as my sister does when studying for a chemistry exam. If we work together, we can finish studying in 2 hours. How long would I take on my own?

OpenStudy (anonymous):

Procedure: 1. Write a word problem describing the time it takes to complete an activity individually and with a friend. (just posted this part) 2. Write a rational equation based upon the word problem you created. 3. Solve the rational equation. Show your work.

OpenStudy (phebe):

@wio

OpenStudy (anonymous):

Okay, I'm trying to do the 2nd part on my own.

OpenStudy (anonymous):

1/x + 1/2x = 1/2

OpenStudy (anonymous):

@wio Help me learn how to solve this?

OpenStudy (anonymous):

I'm going to start by finding an LCD...multiply the first fraction by 2/2, the second stays the same, the third is multiplied by x/x.

OpenStudy (anonymous):

2/2x + 1/2x = x/2x

OpenStudy (anonymous):

2 + 1 = x x = 3 It takes my sister 3 hours to study alone, but because I take twice as long, it takes me 6 hours.

OpenStudy (anonymous):

Is all this okay?

OpenStudy (anonymous):

This?\[ \frac1x + \frac1{2x}=\frac 12 \]

OpenStudy (anonymous):

Yes. And everything past it looks good?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

In general, you can just multiply by the LCD and that is good enough.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

If I have another question in a few minutes, can you stick around?